# PAD: Phase-Amplitude Decoupling Fusion for Multi-Modal Land Cover Classification

### v2

@@ IV-C Experimental Results on Datasets
SAR & RGB MCANet [28] JAG’21 64.21 45.49 42.93 56.20 81.81 25.93 8.47 80.80 45.47 59.94 46.44
CEN [33] TPAMI’22 66.46 56.28 45.46 58.09 82.34 32.51 17.82 82.32 47.47 65.36 51.28
TokenFusion [52] CVPR’22 67.87 57.26 46.96 61.66 83.31 36.34 18.76 83.38 49.87 67.09 53.16
DDHRNet [29] JAG’22 59.12 45.67 30.77 47.96 79.59 0.47 0.03 78.00 27.18 48.36 37.66
CDDFuse [12] CVPR’23 59.88 35.08 25.22 52.31 78.77 17.64 15.00 77.02 38.69 54.29 40.56
CMX [31] TITS’23 69.24 58.20 48.45 64.06 83.83 39.99 21.92 84.18 52.76 69.01 55.10
MRFS [30] CVPR’24 69.39 58.13 49.43 64.90 83.87 41.26 21.72 84.21 54.33 69.38 55.53
FTransUNet [6] TGRS’24 64.87 53.84 43.71 55.63 81.53 0.00 0.00 81.25 43.53 52.97 42.80
ASANet [32] ISPRS’24 69.18 56.95 48.58 64.06 83.79 41.48 21.74 84.15 52.62 69.05 55.11
@@
IV-C1 Evaluations on WHU-OPT-SAR
As shown in Table II, PAD achieves new state-of-the-art results: 56.26% mIoU, 84.56% OA, 53.93% mKappa, and 70.14% mF1. This surpasses the best unimodal method (DANet-RGB [18]: 53.25% mIoU, 83.19% OA) and the multi-modal baseline MRFS [30] (55.53% mIoU, 84.21% OA). In fine-grained classes, PAD attains 43.24% IoU on roads and 49.70% on villages, improving over MRFS by 1.98% and 0.27%, respectively.
Figure 6: Visualization Results on WHU-OPT-SAR. Four representative scenarios: (a) building-dense area, (b) specular water bodies, (c) vegetated region, and (d) urban-rural transition. For each scenario, images are shown (left to right): RGB, SAR, pseudo-color ground truth, and prediction maps from six competing methods. We compare PAD against single-modal baselines (DANet-RGB, TransUNet-SAR) and multi-modal approaches (CDDFuse, CMX, MRFS). Regions of interest are highlighted with gray boxes (buildings/water) and circles (vegetation); see Section IV-C for details.
Visual results in Fig. 6 further illustrate PAD’s benefits. In (a), PAD recovers occluded road segments (cyan) in building-dense areas via SAR phase cues, whereas RGB-only models miss them. In (b), PAD distinguishes specular reflections from true water boundaries under cloud interference. In (c)-(d), PAD resolves vegetation spectral ambiguities and maintains consistency across urban-rural transitions.
TABLE III: Quantitative Results (%) on DDHR-SK. We report per-category IoU for seven land cover classes (Buildings, Roads, Greenery, Water, Farmland), along with OA, mKappa, mF1, and mIoU.
Method Venue IoU per Category OA mKappa mF1 mIoU
@@ IV-C1 Evaluations on WHU-OPT-SAR
SAR & RGB MCANet [28] JAG’21 85.12 47.88 78.69 96.96 73.18 89.59 46.13 85.55 76.37
CEN [33] TPAMI’22 91.78 58.73 87.70 98.33 85.73 94.26 60.19 90.93 84.46
TokenFusion [52] CVPR’22 93.90 70.67 90.62 98.94 89.21 95.79 69.64 93.70 88.67
DDHRNet [29] JAG’22 79.78 13.85 69.85 93.73 62.17 84.73 24.59 73.75 63.88
CDDFuse [12] CVPR’23 96.59 83.02 94.59 99.11 93.30 97.61 80.65 96.46 93.32
CMX [31] TITS’23 96.43 82.59 94.37 99.38 93.72 97.58 81.08 96.44 93.30
MRFS [30] CVPR’24 96.48 83.68 94.48 99.39 93.84 97.65 82.24 96.60 93.57
FTransUNet [6] TGRS’24 84.69 35.21 76.51 96.40 71.11 88.56 36.55 82.35 72.78
ASANet [32] ISPRS’24 95.04 77.09 92.64 99.24 91.82 96.73 75.77 95.21 91.17
@@
IV-C2 Evaluations on DDHR-SK
As shown in Table III, PAD exhibits superior performance under adverse (cloudy) conditions on DDHR-SK. It achieves 94.59% mIoU and 97.65% OA, outperforming leading unimodal methods (e.g., Mask2former-RGB: 90.25% mIoU, 94.76% mF1; DANet-SAR: 83.31% mIoU, 90.28% mF1) and the multi-modal baseline MRFS (93.57% mIoU, 82.24% mKappa). PAD also attains 82.24% mKappa and 96.60% mF1, with boundary-sensitive categories achieving 99.47% IoU on water and 86.12% IoU on roads, surpassing MRFS by 2.44% and 0.08%, respectively. Qualitative results in Fig. 7 further demonstrate PAD’s ability to delineate water infiltration in irrigated fields and maintain continuous road networks beneath fog.
Figure 7: Visualization Results on DDHR-SK. Four boundary-focused scenarios: (a) coastal demarcation, (b) water-cropland interface, (c) vegetation-building complex, and (d) road network topology. For each scenario, panels show the RGB input, SAR coherence map, pseudo-color ground truth, and prediction maps from six methods. We compare PAD against unimodal baselines (Mask2former-RGB, DANet-SAR) and multi-modal approaches (CDDFuse, CMX, MRFS). Critical boundary regions are highlighted with orange boxes; see Section IV-C for details.
The visual results further demonstrate PAD’s capability for cross-modal boundary refinement. In Fig. 7(a), PAD accurately reconstructs periodically submerged coastlines via SAR coherence analysis, whereas RGB-only models confuse tidal sediments with terrestrial areas. In Fig. 7(b), PAD leverages SAR features to delineate water-infiltration boundaries in irrigated fields, overcoming the over-segmentation common in optical sensors. Moreover, in Fig. 7(c) and Fig. 7(d), PAD achieves continuous road networks under vegetation occlusion and precise material boundaries in vegetation-building complexes, respectively.
TABLE IV: Computational Efficiency Comparison. Model size (#Params), inference time per image, and throughput (FPS) for competitive methods. Lower #Params and inference time, and higher FPS, indicate better efficiency.
Method #Params (M) Inference Time (ms) FPS (frames/s)
CMX [31] 15.00 206.70 1,957.8
MRFS [30] 10.30 695.85 1,342.2
ASANet [32] 7.82 72.91 2,245.2
PAD (Ours) 3.88 5.16 3,228.7
IV-C3 Computational Efficiency Analysis
We evaluate the computational efficiency of PAD in terms of model size (Fusion Modules Only), inference time per image, and frames per second (FPS), as summarized in Table IV. We compare against four high-mIoU methods, CMX [31], MRFS [30], ASANet [32], and CDDFuse [12] (excluded from detailed analysis due to its two-stage design and lower efficiency). All metrics are aggregated across fusion stages.
As shown in Table IV, PAD uses only 3.882 M parameters, achieves the fastest inference time (5.16 ms per image), and delivers the highest FPS (3,228.7). These results demonstrate that PAD not only matches or exceeds accuracy of existing methods but also offers real-time efficiency suitable for operational deployment.
IV-C4 Ablation Study
We performed comprehensive ablation studies on WHU-OPT-SAR to validate the contributions of our core components: the Amplitude Spectrum Fusion (ASF) module, the Phase Spectrum Correction (PSC) module, and Multi-Stage Fusion (MSF).
[…]
ID Components IoU per Category OA mKappa mF1 mIoU
MSF ASF PSC Farmland City Village Water Forest Road Others
[…]
@@ IV-C4 Ablation Study
Eight variants (see Table V) were evaluated. The baseline (I), which replaces MSF with early fusion and substitutes PAD (ASF+PSC) with pixel-wise addition, achieves the lowest mIoU (54.98%). Enabling MSF alone (II) yields a +0.39% gain, ASF alone (III) reaches 55.86% mIoU, and PSC alone (IV) provides a smaller improvement. Combining MSF with ASF (V) or PSC (VI) yields similar gains (56.11-56.12% mIoU). Integrating all three modules achieves the best performance (56.26% mIoU), demonstrating their effective synergy.
TABLE VI: Ablation Study on Fusion Modules (%) for WHU-OPT-SAR. We compare alternative amplitude-only, phase-only, and full fusion designs by integrating each into our framework.
Fusion Module IoU per Category OA mKappa mF1 mIoU
@@
Ours 69.91 58.51 49.53 65.52 84.05 42.63 22.58 84.55 54.53 69.93 56.11
Phase FelNFN [16] 60.66 53.25 31.99 54.14 80.64 19.84 15.02 79.42 33.32 58.89 45.08
@@
Ours 69.98 57.32 49.36 65.62 84.12 43.42 23.01 84.57 53.80 69.98 56.12
Whole FelNFN [16] 66.43 51.97 44.36 58.48 82.80 32.33 12.64 82.33 47.35 63.63 49.86
FCENet++ [14] 66.87 53.01 46.41 60.30 82.64 30.21 14.51 82.78 43.83 64.33 50.56
@@
To further validate ASF and PSC, we replaced each with five existing frequency-domain designs, FelNFN [16], FCENet [14], FGSwin [53], Frefusion [15], and SFINet++ [46], while keeping the rest of our framework unchanged (see Table VI). These methods represent common strategies such as fixed-kernel spectral convolution (FelNFN), frequency-spatial interaction via Hadamard products/matrix multiplication (FCENet), learned spectral filtering (FGSwin), attention-based cross-domain fusion (Frefusion), and amplitude-phase parallel processing (SFINet++). Our ASF outperforms other amplitude-specific methods (56.11% vs. 55.85% mIoU for FGSwin), and PSC consistently exceeds phase baselines (56.12% vs. 56.05% for channel attention and 45.08% for FelNFN). At the full-fusion level, our integrated design still leads (56.26% mIoU), indicating both modules are effective in isolation and synergistic when combined. These results confirm that adaptive frequency modulation (ASF) and spatially guided phase correction (PSC) offer structurally unique and more effective alternatives to existing paradigms.
TABLE VII: Ablation Study on Loss Function Weights (%) for WHU-OPT-SAR. We vary $\lambda_{1}$ with fixed $\lambda_{2}=0.1$ and $\lambda_{2}$ with fixed $\lambda_{1}=0.4$.
Configuration IoU per Category OA mKappa mF1 mIoU
@@
We conducted weight ablations (see Table VII) to assess the trade-off between the amplitude consistency loss $\mathcal{L}_{\mathrm{amp}}$ and segmentation loss $\mathcal{L}_{\mathrm{seg}}$. The best mIoU (56.26%) is achieved with $\lambda_{1}=0.4$ and $\lambda_{2}=0.1$. Increasing $\lambda_{1}$ above 0.4 degrades performance, especially on under-represented classes, while too small $\lambda_{2}$ significantly reduces accuracy, confirming segmentation supervision must remain dominant.
TABLE VIII: Ablation Study on Initial Frequency Radius (%) for WHU-OPT-SAR. We vary the trainable radius parameter in the ASF module.
Radius IoU per Category OA mKappa mF1 mIoU
@@
We varied the initial frequency radius $r$, which controls the preserved low-frequency proportion in ASF (see Table VIII). The optimal balance is at $r=0.1$ (56.26% mIoU). Values that overly suppress or over-smooth low-frequency content degrade performance, although the method remains robust for $r\in[0.1,10]$.
IV-D Visualization at Distribution Level
To further analyze feature distributions, we extracted multi-scale features from the RGB and SAR encoders and from the fusion modules. After upsampling to the original resolution, we concatenated features channel-wise. To balance computational efficiency with discriminative power, we applied a two-stage dimensionality reduction: a) random spatial sampling, and b) principal component analysis (PCA), retaining 50 components (over 89% variance). The final visualization was generated using t-SNE [54] (perplexity = 30, 1,000 iterations).

### v3

@@ IV-B Main Quantitative Results
SAR & RGB
MCANet [29] JAG’21 65.14 53.97 45.43 54.59 81.87 31.37 13.43 81.74 44.28 63.37 49.40
DDHRNet [30] JAG’22 64.80 49.42 40.98 54.11 81.31 22.96 10.32 80.94 44.49 59.84 46.27
CEN [34] TPAMI’22 67.60 55.73 46.46 60.65 83.28 34.22 15.94 83.17 49.37 65.79 51.98
TokenFusion [55] CVPR’22 68.43 58.34 45.76 61.09 83.79 32.03 15.20 83.65 49.45 65.67 52.09
CDDFuse [13] CVPR’23 60.07 48.64 23.96 50.67 79.47 15.16 11.43 78.02 51.36 54.54 41.34
CMX [32] TITS’23 70.04 58.41 49.04 64.77 84.12 41.76 22.21 84.53 53.46 69.60 55.76
FTransUNet [6] TGRS’24 68.22 56.28 46.76 62.35 83.46 35.35 17.23 83.50 50.94 66.61 52.81
ASANet [33] ISPRS’24 69.71 58.09 49.05 63.52 84.05 41.50 20.57 84.29 52.83 69.04 55.21
MRFS [31] CVPR’24 69.59 58.04 49.15 63.78 84.15 41.26 22.88 84.24 54.14 69.48 55.55
ASMFNet [56] JSTARS’24 67.31 55.73 46.56 59.51 83.03 32.63 14.60 82.91 49.51 65.09 51.34
MIEFNet [57] GRSL’24 68.72 57.16 48.04 63.23 83.64 39.35 18.41 83.87 50.58 67.89 54.08
MFNet [58] TGRS’25 68.97 58.39 47.78 63.38 83.88 39.51 20.33 84.02 51.86 68.47 54.61
@@
IV-B1 Evaluations on WHU-OPT-SAR
As shown in Table II, PAD achieves new state-of-the-art results: 56.26% mIoU, 84.56% OA, 53.93% mKappa, and 70.14% mF1. This surpasses the best unimodal method (DANet-RGB [19]: 53.25% mIoU, 83.19% OA) and the multi-modal baseline CMX [32] (55.76% mIoU, 84.53% OA). For fine-grained classes, PAD attains 43.24% IoU on roads and 49.70% on villages, improving over MRFS by 1.48% and 0.66%, respectively.
Figure 6: Visualization Results on WHU-OPT-SAR. Four representative scenarios: (a) building-dense area; (b) specular water bodies; (c) vegetated region; and (d) urban-rural transition. For each scenario, images are shown (left to right): RGB, SAR, pseudo-color ground truth, and predictions from six competing methods. We compare PAD against unimodal baselines (DANet-RGB and TransUNet-SAR) and multi-modal approaches (CDDFuse, CMX, and MRFS). Regions of interest are highlighted with gray boxes (buildings/water) and circles (vegetation); see §IV-B for details.
Qualitative results in Fig. 6 further highlight PAD’s benefits. In Fig. 6(a), PAD recovers occluded road segments (cyan) in building-dense areas via SAR phase cues, whereas RGB-only models miss them. In Fig. 6(b), PAD distinguishes specular reflections from true water boundaries under cloud interference. In Fig. 6(c)-(d), PAD resolves vegetation spectral ambiguities and maintains consistency across urban-rural transitions.
TABLE III: Quantitative Results (%) on DDHR-SK. Per-class IoU is reported for five land-cover categories (Buildings, Roads, Greenery, Water, and Farmland), along with OA, mKappa, mF1, and mIoU.
Method Venue Per-class IoU OA mKappa mF1 mIoU
@@ IV-B1 Evaluations on WHU-OPT-SAR
SAR & RGB
MCANet [29] JAG’21 83.90 38.06 75.40 96.04 69.99 88.07 41.04 82.54 72.68
DDHRNet [30] JAG’22 82.76 29.45 73.78 95.20 67.40 87.21 35.24 79.81 69.72
CEN [34] TPAMI’22 86.61 40.08 79.34 97.57 74.26 90.20 46.91 84.50 75.57
TokenFusion [55] CVPR’22 92.28 65.60 88.05 98.60 85.74 94.63 64.71 92.10 86.05
CMX [32] TITS’23 93.50 72.03 90.02 99.00 88.03 95.59 70.48 93.65 88.52
CDDFuse [13] CVPR’23 82.19 35.87 69.95 93.94 54.50 85.06 66.26 78.55 67.29
FTransUNet [6] TGRS’24 86.78 40.51 78.70 97.24 73.13 89.91 46.02 84.35 75.27
ASANet [33] ISPRS’24 86.67 43.18 79.02 97.28 73.99 90.17 46.40 85.03 76.03
MRFS [31] CVPR’24 94.63 76.20 91.30 99.11 90.22 96.32 74.23 94.72 90.29
ASMFNet [56] JSTARS’24 86.24 37.52 79.08 97.45 73.38 89.92 45.24 83.77 74.74
MIEFNet [57] GRSL’24 90.56 63.90 85.20 98.39 82.19 93.46 61.20 90.89 84.05
MFNet [58] TGRS’25 89.79 56.65 83.61 98.10 80.32 92.67 57.54 89.23 81.70
@@
IV-B2 Evaluations on DDHR-SK
As shown in Table III, PAD exhibits superior performance under adverse (cloudy) conditions on DDHR-SK. It achieves 94.59% mIoU and 98.04% OA, outperforming leading unimodal methods (e.g., Mask2Former-RGB: 90.25% mIoU, 94.76% mF1; DANet-SAR: 83.31% mIoU, 90.28% mF1) and the multi-modal baseline MRFS (90.29% mIoU, 74.23% mKappa). PAD also attains 84.38% mKappa and 97.16% mF1. Boundary-sensitive categories achieve 86.12% IoU on roads, surpassing MRFS by 9.92%.
Figure 7: Visualization Results on DDHR-SK. Four boundary-focused scenarios: (a) coastal demarcation; (b) water-cropland interface; (c) vegetation-building complex; and (d) road-network topology. For each scenario, panels (left to right) show the RGB input, SAR coherence map, pseudo-color ground truth, and predictions from six methods. We compare PAD against unimodal baselines (Mask2Former-RGB and DANet-SAR) and multi-modal approaches (CDDFuse, CMX, and MRFS). Critical boundary regions are highlighted with orange boxes; see §IV-B for details.
Visual results demonstrate PAD’s ability for cross-modal boundary refinement. In Fig. 7(a), PAD accurately reconstructs periodically submerged coastlines via SAR coherence analysis, whereas RGB-only models confuse tidal sediments with terrestrial areas. In Fig. 7(b), PAD leverages SAR features to delineate water-infiltration boundaries in irrigated fields, overcoming the over-segmentation common in optical sensors. In Fig. 7(c)-(d), PAD produces continuous road networks under vegetation occlusion and precise material boundaries in vegetation-building complexes.
IV-C Ablation and Efficiency Analysis
IV-C1 Ablation Study
We conduct comprehensive ablations on WHU-OPT-SAR to validate the contributions of our core components: the Amplitude Spectrum Fusion (ASF) module, and the Phase Spectrum Correction (PSC) module.
TABLE IV: Ablation Results (%) on Main Modules for WHU-OPT-SAR.
and
indicate the inclusion or exclusion of ASF and PSC, respectively.
Components Per-class IoU OA mKappa mF1 mIoU
ASF PSC Farmland City Village Water Forest Road Others
69.46 57.16 49.01 65.15 83.87 41.57 21.39 84.32 51.96 69.23 55.37
69.91 58.51 49.53 65.52 84.05 42.63 22.58 84.55 54.53 69.93 56.11
69.98 57.32 49.36 65.62 84.12 43.42 23.01 84.57 53.80 69.98 56.12
69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
@@ IV-C1 Ablation Study
Four variants (see Table IV) are evaluated. The baseline, which replaces substitutes PAD (ASF+PSC) with pixel-wise addition, achieves the lowest mIoU (55.37%). Enabling ASF alone reaches 56.11% mIoU; PSC alone provides 56.12% mIoU. Integrating two modules achieves the best performance (56.26% mIoU), demonstrating effective synergy.
TABLE V: Ablation on Fusion Modules (%) for WHU-OPT-SAR. Amplitude-only, phase-only, and combined (amplitude+phase) fusion designs are compared, each integrated into our framework.
Fusion module Per-class IoU OA mKappa mF1 mIoU
@@
PAD (Ours) 69.91 58.51 49.53 65.52 84.05 42.63 22.58 84.55 54.53 69.93 56.11
Phase
FelNFN [17] 60.66 53.25 31.99 54.14 80.64 19.84 15.02 79.42 33.32 58.89 45.08
@@
PAD (Ours) 69.98 57.32 49.36 65.62 84.12 43.42 23.01 84.57 53.80 69.98 56.12
Whole
FelNFN [17] 66.43 51.97 44.36 58.48 82.80 32.33 12.64 82.33 47.35 63.63 49.86
FCENet [15] 66.87 53.01 46.41 60.30 82.64 30.21 14.51 82.78 43.83 64.33 50.56
@@
To further validate ASF and PSC, we replace each with five existing frequency-domain designs, FeINFN [17], FCENet [15], FGSwin [59], Frefusion [16], and SFINet++ [50], while keeping the rest of our framework unchanged (see Table V). These methods represent strategies such as fixed-kernel spectral convolution (FeINFN), frequency-spatial interaction via Hadamard products/matrix multiplication (FCENet), learned spectral filtering (FGSwin), attention-based cross-domain fusion (Frefusion), and amplitude-phase parallel processing (SFINet++). Our ASF outperforms other amplitude-focused methods (56.11% vs. 55.85% mIoU for FGSwin), and PSC consistently exceeds phase baselines (56.12% vs. 56.05% for channel attention and 45.08% for FeINFN). At the full-fusion level, our integrated design still leads (56.26% mIoU), indicating that both modules are effective in isolation and synergistic when combined. These results confirm that adaptive frequency modulation (ASF) and spatially guided phase correction (PSC) offer structurally distinct and more effective alternatives to existing paradigms.
TABLE VI: Ablation on Loss-Function Weights (%) for WHU-OPT-SAR. We vary $\lambda_{1}$ (with $\lambda_{2}$ fixed at 0.1) and $\lambda_{2}$ (with $\lambda_{1}$ fixed at 0.4).
Configuration Per-class IoU OA mKappa mF1 mIoU
@@
We conduct weight ablations (see Table VI) to assess the trade-off between the amplitude consistency loss $\mathcal{L}_{\mathrm{amp}}$ and the segmentation loss $\mathcal{L}_{\mathrm{seg}}$. The best mIoU (56.26%) is achieved with $\lambda_{1}=0.4$ and $\lambda_{2}=0.1$. Increasing $\lambda_{1}$ above 0.4 degrades performance, especially on under-represented classes, whereas too small $\lambda_{2}$ significantly reduces accuracy, confirming that segmentation supervision must remain dominant.
TABLE VII: Ablation on Initial Frequency Radius (%) for WHU-OPT-SAR. The trainable radius parameter in the ASF module is varied.
Radius Per-class IoU OA mKappa mF1 mIoU
@@
We vary the initial frequency radius $r$, which controls the preserved low-frequency proportion in ASF (see Table VII). The optimal balance is at $r=0.1$ (56.26% mIoU). Values that overly suppress or over-smooth low-frequency content degrade performance, although the method remains robust for $r\in[0.1,10]$.
TABLE VIII: Computational Efficiency Comparison. Model size (#Params), per-image inference time, and throughput (FPS) for competing methods. Fewer parameters and shorter inference times, along with higher FPS, indicate better efficiency.
Method #Params (M) Inference time (ms) Throughput (FPS)
CMX [32] 15.00 206.70 1,957.8
MRFS [31] 10.30 695.85 1,342.2
ASANet [33] 7.82 72.91 2,245.2
PAD (Ours) 3.88 5.16 3,228.7
IV-C2 Computational Efficiency
We evaluate computational efficiency in terms of model size (fusion modules only), inference time per image, and frames per second (FPS), summarized in Table VIII. We compare against three high-mIoU methods, CMX [32], MRFS [31], and ASANet [33]. All metrics are aggregated across fusion stages with the same baseline model.
As shown in Table VIII, PAD uses only 3.882 M parameters, achieves the fastest inference time (5.16 ms per image), and delivers the highest FPS (3,228.7). These results indicate that PAD not only matches or exceeds the accuracy of existing methods but also offers real-time efficiency suitable for operational deployment.
IV-D Further Analysis and Visualization
IV-D1 Visualization at the Distribution Level
To further analyze feature distributions, we extract multi-scale features from the RGB and SAR encoders and from the fusion modules. After upsampling to the original resolution, we concatenate features channel-wise. To balance computational efficiency with discriminative power, we apply a two-stage dimensionality reduction: a) random spatial sampling and b) principal component analysis (PCA), retaining 50 components (over 89% variance). The final visualization is generated using t-SNE [60] (perplexity = 30, 1,000 iterations).